# Deciphering the global production network from cross-border firm transactions

### v2

@@ Abstract
Critical for policy-making and business operations, the study of global supply chains has been severely hampered by a lack of detailed data. Here we harness global firm-level transaction data covering 20m global firms, and 1 billion cross-border transactions, to infer key inputs for over 1200 products. Transforming this data to a directed network, we find that products are clustered into three large groups including textiles, chemicals and food, and machinery and metals. European industrial nations and China dominate critical intermediate products in the network such as metals, common components and tools, while industrial complexity is correlated with embeddedness in densely connected supply chains. To validate the network, we find structural similarities with two alternative product networks, one generated via LLM queries and the other derived by NAFTA to track product origins. We further detect linkages between products identified in manually mapped single sector supply chains, including electric vehicle batteries and semi-conductors. Finally, metrics derived from network structure capturing both forward and backward linkages are able to predict country-product diversification patterns with high accuracy.
Keywords: Supply chains, resilience, policy, network science
Disclaimer: This paper contains the views of the authors and not necessarily those of the UK Government or its Ministers.
Acknowledgements: We would like to thank Christian Diem and Tobias Reisch for their useful input and comments, and Shaun Hoang for his research assistance. The data and computing infrastructure was provided by the the UK Government Global Supply Chain Intelligence Programme (GSCIP). NOC would like to acknowledge funding from the 2024-2025 UKRI Policy Fellowship programme.
@@ Introduction
As acutely highlighted by the global pandemic, limited insight into global supply chains can have serious and immediate consequences. However, until quite recently, supply chain analysis was limited to specific sectors or highly aggregate economy-wide transaction data which lacks detail and hence is not very useful or sufficiently accurate for many policy and research questions. Here we harness ’next generation’ global inter-firm transaction data to infer detailed product inputs as the basis for constructing a product level supply chain network.
@@
Traditionally, flows and transactions between sectors in the economy have been captured via input-output tables (Leontief, 1986). These capture monetary flows between firms within defined industries, and are typically produced by national statistical agencies. Some countries, such as the US and Mexico, produce these at quite a detailed level (4 digit NAICS - or approx 400 industries). However, in most countries these are only available for very aggregate sectors (in the UK, for example, IO tables cover just 105 sectors). And, importantly, they don’t capture cross border flows limiting their potential for global supply chain analysis. There have been a handful of efforts to build integrated input-output networks that include inter-country flows, but again these are extremely aggregate in nature. For example, the World Input-Output Dataset (WIOD) (Timmer et al., 2012) has just 56 sectors and 40 countries (2010-14) or the OECD World Input-Output Table (WIOT) with 36 sectors and 69 countries (1995-2020). A variety of studies have investigated the structure of the World Input-Output Network (WION), finding strong geographical clustering of sectors within neighbouring countries, but also some instances of strong inter-country clustering in sectors such as textiles and transport equipment Piccardi et al. (2018).
In recent years there has been a huge push to better collect and make available for research firm level transaction data (Pichler et al., 2023). Research using this data has indicated that, for example, the propagation of shocks at the firm level has distinct characteristics and magnitudes which cannot be approximated by aggregate data (Diem et al., 2022, 2024; Inoue and Todo, 2024). Up until very recently, however, most firm level data was collected by single countries (e.g. VAT data) or for a specific purpose (i.e., shipping data), and hence had very limited coverage.
@@
At the other end of the spectrum are efforts to manually reconstruct detailed supply chains for specific sectors. These studies collect data from a variety of sources including internal systems like procurement, inventory management, and logistics, as well as external sources such as supplier data and transportation information, to build a detailed picture of the key inputs needed to produce often strategically important goods. For example, there have been careful efforts to map out antibiotic (Klimek et al., 2023), semi-conductor (Haramboure et al., 2023) and battery supply chains (Caravella et al., 2024). However, these supply chain mapping exercises tend to focus on material inputs - critical minerals and chemicals - rather than machinery, tools and other necessary inputs such as cleaning equipment. And, critically, their fragmented nature makes them not suited to many questions which require a wider structural lens on product inter-connectivity.
@@
A major focus of analytical efforts with respect to supply chains has been to been to better understand complex global inter-dependence structures and vulnerability to shocks. Risks might come from geographic shocks, such as localised shipping disruptions or natural disasters, sector specific shocks, perhaps due to tariffs or missing critical inputs, or broad based shocks such as the pandemic or financial crisis. Much of this literature aims to quantify node importance - whether it be nodes that influence the degree of spreading processes, or those that are most susceptible to impacts. A number of papers suggest that the US IO network follows a power law structure (Acemoglu et al., 2012; Han and Goetz, 2019; Chopra and Khanna, 2015), and that a small number of key sectors (including critical infrastructure sectors) are involved in many economic transactions and are hence susceptible to inducing large scale cascades. Blöchl et al. (2011) proposed two centrality measures to capture the impact of shocks: random walk centrality to capture most immediate impacts, and betweenness to capture where a shock lingers longest. The authors apply these metrics to the OECD WIOT, finding ’wholesale and retail trade’ is most often the node with highest random walk centrality while ’motor vehicles’ have the highest betweenness.
Supply chain linkages are also key to longer term impacts such as diversification of countries and regions into new economic activities, otherwise referred to as ’upgrading’ (Kaplinsky, 2000). Originally proposed as a mechanism to aid countries with natural resources diversify into productive activities (Hirschman, 1958; Gereffi et al., 1999), it is widely understood today that embeddedness in global supply chains aids both domestic productivity and export diversification (Pahl and Timmer, 2020). While a wide array of mechanisms have been proposed, many focus on accumulation of technological capabilities and the diffusion of knowhow, as espoused by the Evolutionary Economic Geography (Nelson and Winter, 2002; Boschma and Frenken, 2006) and Innovation Systems (Edquist, 2013) literatures. From this perspective, supply chain linkages are thought to drive knowledge diffusion via, for example, demands for quality upgrading, worker flows and increased competition (Blomström and Kokko, 1998; Criscuolo et al., 2017). These effects are thought to be particularly strong in the case of supplier-customer relationships between local firms and co-located foreign multi-nationals (Javorcik, 2004), but there are greater uncertainties about the potential for such linkages to induce upgrading when it comes to linkages spanning large geographical and technological distances (Pietrobelli and Rabellotti, 2011).
The empirical evidence on the benefits of participation in global value chains is mixed (as reviewed by Kowalski et al. (2015)), finding a strong influence on manufacturing productivity and export sophistication, but mixed impact, for example, on domestic employment (Pahl and Timmer, 2020; Gonzalez, 2016). In terms of direction, both backward linkages (i.e., diversification into more complex downstream goods) and forward linkages (i.e., into input or supplier goods) have been proposed as channels for diversification. Methodologically closest to our work, Bahar et al. (2019) find evidence for the role of forward linkages in international export product diversification patterns over a 10 year period. Linkages in this case, however, were proxied from US IO data measured for US domestic industrial sectors and do not capture global value chain relations.
@@
Beyond the short-medium term impacts of shocks, a recent literature studies the influence of production network structure on local economic growth and resilience. Taking an evolutionary stance, in this context resilience is typically defined as the extent to which regions or countries adapt to shocks in the longer term rather than return to pre-shock norms (Boschma, 2015). Empirical evidence at a subnational level supports the idea that regions’ diversification patterns are conditioned on their ’role’ in national production networks. For example, US cities tend to enter new industries that are ’related’ to those that already exist (Essletzbichler, 2017), where relatedness is measured in terms of supply chain proximity (Fan and Lang, 2000). When it comes to resilience, Han and Goetz (2019) show that weighted centrality metrics derived from the US IO network are associated with employment resilience for US counties during the 2007-2009 recession, while Cainelli et al. (2019) find that EU regions with highly integrated production networks were worse off during the same period (which the authors suggest is due to enhanced amplification of the shock).
@@
Here we aim to infer detailed ’product recipes’, or the set of unique inputs needed in the production of a specific product. In other words, we aim to capture the input materials and tools used in product that are specific to a particular product, and not common across many products (e.g., packaging). Distinct from two recent efforts build a similar product network using firm-level transaction data from Italy (Fessina et al., 2025) and Japan (Inoue and Todo, 2024), we aim to capture global supply chains (as opposed to local supply chains) via global firm purchasing patterns. To do this, we harness ’next generation’ global firm-level transaction data via a UK government programme, GSCIP. We use an aggregated subset of data from GSCIP for the years 2021-2023, prior to additional augmentation of the transactional data completed in 2025. The subset of data that we use is orders of magnitude larger than single country datasets typically in use today, and covers 20m global firms, and 1 billion cross-border transactions. The data was assembled from a variety of sources, from customs to firm-owned data, and matched at a firm level.
@@
To detect the set of inputs required to make a specific product, we compute the excess purchase of inputs (products) for the set of firms that make a specific product relative to how often an average firm purchases the input. This latter step effectively down-weights common or ubiquitous inputs. This approach echoes that of O’Clery et al. (2021), in which the authors inferred country-product diversification paths by extracting the set of common products produced in the set of countries in which a specific new product appeared later. In this case, one can think of this approach as ’piecemeal’ reconstruction of global product supply chains in that we independently estimate each set of input linkages for each product. In this way, we capture both inputs or ingredients, and complex tools and packaging.
@@
These relationships can be represented and analysed as a network in which nodes are products, and directed weighted edges represent the relative importance of the input. Analysis of the meso-structure of this network suggests that textiles’ supply chains are largely isolated from wider manufacturing, while chemicals are typically more inter-connected with food supply chains than machinery and metals. European industrial nations and China dominate nodes with high betweenness centrality, i.e., critical intermediate products in the network, while countries with high industrial complexity (proxied via the ECI metric of Hidalgo and Hausmann (2009)) tend to export goods that have a high ’hub score’ in the network, i.e., they are embedded in densely connected supply chains.
@@
While there is no ‘ground truth’ data available, we have undertaken significant efforts to validate the network structure. We found structural similarities with an alternative AI-generated product network (Fetzer et al., 2024), and a product supply chain dataset created by NAFTA to track product origins (Conconi et al., 2018). In addition, we found dense connectivity in our network between nodes identified as belonging to two manually constructed supply chains - electric vehicle batteries and semi-conductors. Finally, metrics derived from the network structure are able to predict country-product diversification patterns with high accuracy, finding evidence for both forward and backward linkages.
@@ Results
We use global inter-firm transaction data covering over 20m firms and 1 billion transactions. Firm level transaction data has been compiled from a variety of sources including customs and firm-held data, such as shipping data. These sources have been integrated using AI-based tools to, for example, disambiguate firm names and match them. The data we used represents a static cut which was made available to the research team via a UK government programme (GSCIP), includes firm level attributes for over 20m firms and 1 billion inter-firm transactions between 2021-23. All transactions are between firms located in different countries. The GSCIP data contains additional firm-level information from other sources, such as firm ownership structures.
We have a number of firm and transaction attributes, a selection of which is shown in Figure 1 B, including dollar transactions for specific products between firm pairs. As shown in Figure 1 C-D, the data has high coverage across developed economies, and those with high trade (exports plus imports) volumes such as China and Vietnam. It also has good coverage in some middle income countries such as India and Peru, but with lower coverage overall in less economically active regions such as Africa.
Figure 1: We use global inter-firm transaction data covering over 20m firms and 1 billion transactions. [A] Firm level transaction data has been compiled from a variety of sources. [B] This static cut of GSCIP data includes firm level attributes for 20m firms, combined with 1 billion inter-firm transactions. [C-D] The data has very good coverage of firms in developed economies, with lower coverage in less economically active regions such as Africa. Top 50 countries by firm count shown.
@@
More precisely, we set stringent definitions to identify firms that both produce and purchase specific products:
@@
If $A_{i,j}>1$, then product $i$ is more prevalent in the purchasing basket of firms that produce $j$ than the wider set of all firms. We can also set a minimum for the size of the set $|S^{i}_{j}|$ (the number of firms that both produce $j$ and buy $i$), which we refer to as parameter ’firmcount’. In essence, for low values of firmcount, we have few ’pieces of evidence’ for the edge weight (and so for some applications we might want to set a higher firmcount threshold, which would make our network sparser). In the analysis below, we often show results for a range of values of both the edge weight threshold and firmcount.
@@
Figure 2: We infer the importance of each product as an input to any other product from the firm transaction data. [A] We compute the excess purchase of product $i$ by firms that produce product $j$, relative to the purchase of $i$ by ’an average firm’. In our toy example here we show that while aluminium panels are purchased very frequently by washing machine firms relative to all firms, cardboard boxes are purchases by all types of firms including washing machine firms and hence they are a very ubiquitous input (and are thus down-weighted in our scheme). [B] We pick up both inputs or ingredients and complex tools. For example, for tomatoes, we capture everything from wooden stakes to fertilizer, harvesting machines and packaging.
@@
We transform the product input relations into a network, and find three main clusters representing machinery and metals, chemicals and food, and textiles. We transform the product input relations into a directed weighted network, with edge weight stored in adjacency matrix $A$ (with entries $A_{i,j}$ as above) of size $n$ x $n$ for $n=1228$. In Figure 3 A, we observe dense edges within sectors (HS sections), shown in coloured blocks on the diagonal. As we increase the firmcount parameter, the edges become more sparse, but the edge weight threshold has no major effect. In practice, the majority of the analysis to follow focuses on the largest weakly connected component of the network with edge threshold value 2 and firmcount value 2, with size $n=1207$. In many cases, and also in the SI, we also show results when varying these parameters.
@@
This network can be visualised using a ’spring’ based visualisation algorithm (Force Atlas in Gephi was used here), as shown in Figure 3 B. We observe loose clustering of nodes according to HS ’Section’, a high level classification of products into sectors. However, rather than rely on an external classification which bears no relation to the inter-connectivity we observe in our data, we use a multi-scale community detection algorithm to uncover the ’meso’ structure or high-level organisation of the network into node ’communities’. These are clusters of products with strong internal connectivity - but relatively weaker connectivity to the rest of the network. These structural patterns are highly informative, capturing for example the behaviour of cascades and shock propagation (Fortunato, 2010).
While there are many distinct approaches to community detection suited to different applications (Fortunato, 2010), in our case, we employ a multi-scale algorithm based on random walkers on a network (Delvenne et al., 2010). Essentially, as the walkers roam on the network, jumping from node to node based on probabilities derived from the directed edge weights, we detect larger and larger clusters with strong internal connectivity. Hence, as ’time’ goes on, the algorithm detects fewer but larger communities. We can learn from this process: as the algorithm effectively ’forces’ communities to merge into larger clusters, the ’time’ at which they merge (or refuse to merge) is highly informative about their relative inter-connectivity. In other words, those that merge earlier are more connected, but those that remain unmerged for longer times are more isolated and disconnected from the wider network.
Here, shown in Figure 3 C-E, we show three network partitions into communities at three distinct ’times’ (see SI for more details and further analysis). The merging process suggests that textiles is highly isolated in the sense that it remains unmerged across all three partitions shown. We find that initially (Figure 3 C) machinery is split across two clusters - one focused on electrical goods and the other on metal-based tools (and metals themselves). We observe that over time, food merges with plastics and then chemicals, and both machinery clusters merge. Then, these two groupings merge - but textiles and some other small clusters (musical products, watches, and art) remain unmerged signifying their isolation. Overall, we find that the product network is broadly segregated into three main clusters representing machinery and metals, chemicals and food, and textiles.
@@
Figure 3: We transform the product input relations into a network. [A] The product supply chain network can represented by an adjacency matrix in which a value in entry $A_{i,j}$ denotes a non-zero edge weight between node (product) $i$ and node $j$. We observe dense edges within sectors (HS sections), shown in coloured blocks on the diagonal. As we increase the firm-count parameter, the edges become more sparse, but the edge weight threshold has no major effect. [B] The network can also be visualised using a visualisation algorithm (Force Atlas was used here). [C-D] We use a multi-scale community detection algorithm to uncover the ’meso’ structure of the network. As the algorithm effectively ’forces’ communities to merge into larger clusters, we find that textiles is highly isolated in the sense that it remains unmerged over many scales. Overall, we find that the product network is broadly segregated into three main clusters representing machinery and metals, chemicals and food, and textiles.
European industrial nations and China dominate critical intermediate products, while industrial complexity is correlated with embeddedness in densely connected supply chains. In the SI we show that machinery products including those related to cars and motors tend to have a large in-degree (they require many inputs), while components, plastics and packaging tend to be inputs to many products. Further, consistent with what we might expect, deploying a UN classification that sorts products into capital, intermediate and consumption goods, we find that capital products are inputs to many products (high out-degree), while both capital and consumption goods have high in-degree – they require many ingredients.
Betweenness centrality (BC) measures the extent to which a node acts as a key ’juncture’ through which many paths (i.e., supply chains) cross. Mathematically, the betweenness centrality of a node is the number of times a path (between any pair of nodes) paths transits that node. In Figure 4 A, we highlight nodes with high betweenness centrality, including various types of machinery, common components such as pumps, blades, taps and motors, plastics, packaging, and chemical binders. Nations which produce these key inputs have an important role in global supply chains, controlling so-called ’choke points’. To investigate this, we compute the Revealed Comparative Advantage (a measure of export concentration of a country in a product) of 113 countries from global trade data (Comtrade 2024). In Figure 4 B, we show a global map with countries shaded by their share of top-50 BC products exported with RCA. We find European industrial nations such as Germany, Italy and Austria, as well as China, dominate high BC products.
Highly industrialised nations engage in complex production processes via embeddedness global supply chains. To capture production complexity, we adopt an approach proposed by Hidalgo and Hausmann (2009) to infer product complexity (PCI) from patterns of geographical production. To measure the complexity of the supply chain in the vicinity of a product in our network, we compute the hub score of each node. This is similar to eigencentrality in the sense that it captures ’neighbour of neighbour’ linkages, but is suitable for directed networks such as ours. We find that the PCI (obtained for 2023 from the Harvard Atlas of Economic Complexity) has a rank correlation of 0.46 with node hub score. In essence, complex products are embedded in complex supply chains. At the country level, the country-level Economic Complexity Index (ECI), which is mathematically equivalent to the mean PCI of products exported with RCA, has a high rank correlation of 0.85 with mean hub score. Hence, countries with high industrial complexity tend to export products which are highly embedded in densely connected global supply chains.

### v3

@@ Abstract
Critical for policy-making and business operations, the study of global supply chains has been severely hampered by a lack of detailed data. Here we harness international firm-level transaction data covering 20m global firms, and 1 billion cross-border transactions, to infer key inputs for over 1200 products. Transforming this data to a directed network, we find that products are clustered into three large groups including textiles, chemicals and food, and machinery and metals. European industrial nations and China dominate critical intermediate products such as metals, common components and tools, while industrial complexity is highly correlated with embeddedness in densely connected supply chains. We find structural similarities with AIPNET, a product network generated via LLM queries, and strong linkages between products identified in manually-mapped electric vehicle battery and semiconductor supply chains. Finally, both forward and backward linkages are predictive of country-product diversification patterns, with stronger overall evidence for backward (upstream) linkages.
Teaser: Harnessing 1 billion cross-border inter-firm transactions to reconstruct the global supply chain production network.
Disclaimer: Any views expressed are solely those of the author(s) and so cannot be taken to represent those of the Foreign Commonwealth and Development Office (FCDO), or to state FCDO Policy. This paper should therefore not be reported as representing the views of the FCDO, its Ministers or His Majesty’s Government.
@@ Introduction
As acutely highlighted by the global pandemic, limited insight into global supply chain dynamics can have serious and immediate consequences. However, until quite recently, supply chain analysis was limited to specific sectors or highly aggregate economy-wide transaction data which lacks detail and hence is not very useful or sufficiently accurate for many policy and research questions. Here we harness ’next generation’ global inter-firm transaction data to infer detailed product inputs as the basis for constructing a product-level global supply chain network.
@@
Traditionally, flows and transactions between sectors in the economy have been captured via input-output (IO) tables [1]. These capture monetary flows between firms within defined industries, and are typically produced by national statistical agencies. Some countries, such as the US and Mexico, produce these at quite a detailed level (4 digit NAICS - or approx 400 industries). However, in most countries they are only available for very aggregate sectors (in the UK, for example, IO tables cover just 105 sectors). And, importantly, they don’t capture cross border flows limiting their potential for global supply chain analysis. There have been a handful of efforts to build integrated input-output networks that include inter-country flows, but again these are extremely aggregate in nature. For example, the World Input-Output Dataset (WIOD) [2] has just 56 sectors and 40 countries (2010-14) and the OECD World Input-Output Table (WIOT) has just 36 sectors and 69 countries (1995-2020).
In recent years there has been a huge push to better collect (and make available for research) firm-level transaction data which has until now been mainly available for individual countries [3]. Research has found common features across these national datasets [4], and that systemic risk in production networks is concentrated in firms that occupy critical positions in product-level supply chains and provide essential intermediate goods [5, 6, 7]. This data is typically, however, collected by single countries (e.g. VAT data) or for a specific purpose (i.e., shipping data), and hence has very limited coverage when it comes to studying global supply chains.
@@
At the other end of the spectrum are efforts to manually reconstruct detailed supply chains for specific sectors. These studies collect data from a variety of sources including internal systems like procurement, inventory management, and logistics, as well as external sources such as supplier data and transportation information, to build a detailed picture of the key inputs needed to produce often strategically important goods. For example, there have been careful efforts to map out antibiotic [8], semi-conductor [9] and critical minerals in electric vehicle battery [10] and defence [11] supply chains. However, these mapping efforts tend to focus on material inputs such as critical minerals and chemicals rather than machinery, tools and other necessary inputs such as cleaning equipment. And, critically, their fragmented nature makes them not suited to many questions which require a wider structural lens on product inter-connectivity.
@@
A major focus of analytical efforts with respect to supply chains has been to been to better understand complex global inter-dependence structures and vulnerability to shocks [12]. Risks might come from geographic shocks, such as localised shipping disruptions or natural disasters [13, 14], sector specific shocks, perhaps due to tariffs or missing critical inputs, or broad-based shocks such as the pandemic [15] or financial crisis. Much of this literature aims to quantify node importance - whether it be nodes that influence the degree of spreading processes, or those that are most susceptible to impacts. For example, a number of papers suggest that the US sectoral IO network follows a power law structure [16, 17, 18], and that a small number of key sectors (including critical infrastructure sectors) are involved in many economic transactions and are hence susceptible to inducing large scale cascades. To improve resilience, some recent work focuses on mitigation strategies such as re-routing [19].
@@
Beyond the short-medium term impacts of shocks, a recent literature studies the influence of production network structure on local economic growth and resilience. Taking an evolutionary stance [20], in this context resilience is typically defined as the extent to which regions or countries adapt to shocks in the longer term rather than return to pre-shock norms [21]. For example, [17] show that US regions with more central industries in the US IO network experience better employment resilience during the 2007-2009 recession, while [22] find that EU regions with highly localised and integrated production networks were worse off during the same period (which the authors suggest is due to enhanced amplification of the shock).
One of the key mechanisms by which supply chains support local economic growth is through sectoral diversification. A key pillar of industrial policy, particularly in resource-rich countries such as African nations with critical minerals, industrial ’upgrading’ is seen as a way for countries to diversify into higher value-added activities [23, 24, 25]. While a wide array of mechanisms have been proposed, many focus on diffusion of knowhow via, for example, demands for quality upgrading, worker flows and increased competition [26, 27]. While these effects have been long documented in the case of supplier-customer relationships between local firms and co-located foreign multi-nationals [28], there are greater uncertainties about the potential for such linkages to induce upgrading when spanning large geographical and technological distances [29].
Known as backward diversification, countries also move into ’upstream’ goods such as machinery or components. Hirschman (1958) [23] suggested that the combination of a local market for the good and know-how on how it works would drive this type of diversification, particularly in developing countries. Upstream goods include both basic inputs and highly complex and high-value goods such as chemicals, electronics and equipment produced by advanced manufacturing economies [30, 31]. Various empirical studies find evidence of upstream diversification [32, 33]. Methodologically closest to our work, Bahar et al (2019) [32] focuses on international export product diversification patterns over a 10 year period, but linkages in this case were proxied from US IO data measured for US domestic industrial sectors and do not capture global value chain relations.
@@
Here we aim to infer detailed ’product recipes’, or the set of unique inputs needed in the production of a specific product. In other words, we aim to capture the input materials and tools that are specific to a particular product (and hence not common across many products, e.g., packaging). Distinct from two recent efforts build a similar product network using firm-level transaction data from specific countries (Italy [34] and Japan [7]), we aim to capture global supply chains. To do this, we harness ’next generation’ global inter-firm transaction data via a UK government programme, GSCIP. The data, procured under license by GSCIP, was compiled by a commercial entity from a variety of sources, from customs to firm-owned data, and matched at a firm level.
@@
To detect the set of inputs required to make a specific product, we compute the excess purchase of each input (product) by the set of firms that make that product relative to how often an average firm purchases the input. This latter step effectively down-weights common or ubiquitous inputs. This approach echoes that of [35] in which the authors inferred typical product diversification paths by extracting the set of common products produced previously by countries which went on later to make a specific product. One can think of our approach here as ’piecemeal’ reconstruction of global product supply chains in that we independently estimate the set of input linkages for each product. In this way, we capture both inputs or ingredients, and complex tools and packaging.
@@
These relationships can be represented and analysed as a network in which nodes are products, and directed weighted edges represent the relative importance of the input. Analysis of the meso-structure of this network suggests that textiles’ supply chains are largely isolated from wider manufacturing, while chemicals are typically more inter-connected with food supply chains than machinery and metals. European industrial nations and China dominate nodes with high betweenness centrality, i.e., critical intermediate products in the network, while countries with high industrial complexity (proxied via the ECI metric of [36]) tend to export goods that have a high ’hub score’ in the network, i.e., they are embedded in densely connected supply chains.
@@
While there is no ‘ground truth’ data available, we have undertaken significant efforts to compare the network structure to other related efforts. We found structural similarities with an alternative AI-generated product network [37], and a product supply chain dataset created by NAFTA to track product origins [38]. In addition, we find dense connectivity in our network between nodes identified as belonging to two manually constructed supply chains - electric vehicle batteries [10] and semiconductors [9]. Finally, metrics derived from the network structure are able to predict country-product diversification patterns with high accuracy. We find evidence for both downstream and upstream linkages, but with overall stronger results for diversification into upstream goods.
@@ Results
We use global inter-firm transaction data covering over 20m firms and 1 billion transactions to construct the product network. Firm level transaction data has been compiled from a variety of sources including customs and firm-held data, such as bill of lading (shipping) data. These sources have been integrated by a commercial provider using AI-based tools to, for example, disambiguate firm names and match them. The data covers transactions between firms located in different countries during the period 2021-23 and was made available to the research team via a UK government programme (GSCIP). GSCIP procured the firm transaction data from a commercial entity, and linked it to additional firm-level information from other sources, such as firm ownership structures (see SI for more detail).
We have a number of firm and transaction attributes, a selection of which are shown in Figure 1 B, including dollar transactions for specific products between firm pairs. While our dataset does not include primary-source data for every country, as shown in Figure 1 C-D, the data has high coverage across developed economies including G7 countries, and those with high trade (exports plus imports) volumes such as China and Vietnam. It also has good coverage in some middle income countries such as India and Peru, but with lower numbers of firms overall in less economically active regions such as Africa. We provide additional information on coverage in the SI.
Figure 1: We use global inter-firm transaction data covering over 20m firms and 1 billion transactions. [A] Firm level transaction data has been compiled from a variety of sources. [B] This includes firm level attributes for 20m firms, combined with 1 billion inter-firm transactions. [C-D] The data has very good coverage of firms in developed economies, with lower coverage in less economically active regions such as Africa. Top 50 countries by firm count shown.
@@
More precisely, we set stringent conditions to identify firms that both produce and purchase specific products:
@@
If $A_{i,j}>1$, then product $i$ is more prevalent in the purchasing basket of firms that produce $j$ than the wider set of all firms. We can also set a minimum for the size of the set $|S^{i}_{j}|$ (the number of firms that both produce $j$ and buy $i$), which we refer to as parameter ’firmcount’. For low values of firmcount, we have few ’pieces of evidence’ for the edge weight - and so for some applications we might want to set a higher firmcount threshold. This would make our network sparser, but we would have higher confidence in the edge weights as they are based on a higher number of signals from the data. In the analysis below, we typically show results for a range of thresholds on both the edge weight ($A_{i,j}$) and firmcount ($|S^{i}_{j}|$).
@@
Figure 2: We infer the importance of each product as an input to any other product from the firm transaction data. [A] We compute the excess purchase of product $i$ by firms that produce product $j$, relative to the purchase of $i$ by ’an average firm’. In our toy example here we show that while aluminium panels are purchased very frequently by washing machine firms relative to all firms, cardboard boxes are purchases by all types of firms including washing machine firms and hence they are a very ubiquitous input (and are thus down-weighted in our scheme). [B] We pick up both inputs or ingredients and complex tools. For example, for beer, we capture everything from hops and malt to containers and bottle stoppers.
@@
We transform the product input relations into a network, and find three main clusters representing machinery and metals, chemicals and food, and textiles. We transform the product input relations into a directed weighted network, with edge weight stored in adjacency matrix $A$ (with entries $A_{i,j}$ as above) of size $n$ x $n$ for $n=1228$ products. In Figure 3 A, we observe dense edges within sectors (HS sections), shown in coloured blocks on the diagonal. As we increase the firmcount parameter, the edges become more sparse, but the edge weight threshold is less sensitive. Unless otherwise stated, the analysis to follow focuses on the largest weakly connected component of the network with edge threshold value 2 and firmcount value 2, with size $n=1207$.
@@
This network can be visualised using a ’spring’ based visualisation algorithm (Force Atlas in Gephi was used here), as shown in Figure 3 B. We observe loose clustering of nodes according to HS ’Section’, a high level classification of products into sectors, but with strong connections between sectors. We use a multi-scale community detection algorithm [39] to uncover the ’meso’ structure or high-level organisation of the network into node ’communities’. These are clusters of products with strong internal connectivity - but relatively weaker connectivity to the rest of the network - and can be seen as an alternative classification of products based on supply chain connectivity. These structural patterns are highly informative, capturing for example the behaviour of cascades and shock propagation [40].
@@
Figure 3: We transform the product input relations into a network. [A] The product supply chain network can represented by an adjacency matrix in which a value in entry $A_{i,j}$ denotes a non-zero edge weight between node (product) $i$ and node $j$. We observe dense edges within sectors (HS sections), shown in coloured blocks on the diagonal. As we increase the firm-count parameter, the edges become more sparse, but the edge weight threshold has no major effect. [B] The network can also be visualised using a visualisation algorithm (Force Atlas was used here). [C-D] We use a multi-scale community detection algorithm [39] to uncover the ’meso’ structure of the network. As the algorithm effectively ’forces’ communities to merge into larger clusters, we find that textiles is highly isolated in the sense that it remains unmerged over many scales. Overall, we find that the product network is broadly segregated into three main clusters representing machinery and metals, chemicals and food, and textiles.
While there are many distinct approaches to community detection suited to different applications [40], in our case, we employ a multi-scale algorithm based on random walkers on a network [39]. As walkers roam on the network, jumping from node to node based on probabilities derived from the directed edge weights, the algorithm detects larger and larger clusters. Hence, as ’time’ increases, the algorithm typically detects fewer but larger communities. We can learn from this process: as the algorithm effectively ’forces’ communities to merge into larger clusters (although this process is not strictly hierarchical), the ’time’ at which they merge (or refuse to merge) is highly informative about their relative inter-connectivity. In other words, those that merge earlier are more connected, but those that remain unmerged for longer times are more isolated and disconnected from the wider network.
Here, in Figure 3 C-E, we show three node partitions corresponding to three distinct ’times’ or scales (see SI Figure S3 for more details and further analysis). The merging process suggests that textiles is highly isolated in the sense that it remains unmerged across all three partitions shown. We find that initially (Figure 3 C) machinery is split across two clusters - one focused on electrical goods and the other on metal-based tools (and metals themselves). As the algorithm detects larger clusters over longer times, food merges with plastics and then chemicals, and both machinery clusters merge. Then, these two groupings merge - but textiles and some other small clusters (musical products, watches, and art) remain unmerged and are thus highly isolated with weak connectivity to the wider network. Overall, we find that the product network is broadly segregated into three main clusters representing machinery and metals, chemicals and food, and textiles.